A certain number is 9 less than another number. The sum of double the great number and 6 is equal to the product of 4 and the smaller number. Select the true statements. 1. The greater number is 21 2. The greater number is 17 3. the smaller number is 8 4. The smaller number is 12.
Answer by sudhanshu_kmr(1152)


x and x-9
2x+ 6 = 4 (x-9)
=> 2x + 6 = 4x - 36
=> 2x= 42
=> x= 21
x-9 = 12
true statements..
1. The greater number is 21
4. The smaller number is 12.